-5(4n+2)=5 i am stuck on this help appreciated
Answer by macston(5194)

.
-5(4n+2)=5 . Divide each side by -5.
4n+2=-1 . Subtract 2 from each side.
4n=-3 . Divide each side by 4.
n=-3/4
.